OpenStudy (anonymous):

Which number equals sqrt(1,029) 7 sqrt(21) 14 sqrt(3) 3 sqrt(21) 21 sqrt(3)

OpenStudy (anonymous):

it is 7sqrt21

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

\[\sqrt{1029}=\sqrt{7*7*21}=7\sqrt{21}\]